OpenStudy (anonymous):

If a twelve or an eight occurs on the first roll, the player wins. What is the probability of winning on the first roll?

OpenStudy (ddcamp):

What sort of die?

OpenStudy (anonymous):

how many sided die?

OpenStudy (anonymous):

or r there 2 dice

OpenStudy (ddcamp):

Assuming that it's 2×6 sided dice: P(12) = 1/36 (6&6) P(8) = 5/36 (2&6, 3&5, 4&4, 5&3, 6&2) P(A or B) = P(A) + P(B)

OpenStudy (anonymous):

and if its a 12 sided die the P(12)=1/12 and P(8)=1/12